OpenStudy (anonymous):

Write a system of equations for the following problems and solve: An errand boy went to the bank to deposit some bills for his employer. Some were one-dollar bills and the rest were five dollar bills. The total value of the bills was $86 and the number of bills was 38. Find the number of each kind of bill.

OpenStudy (anonymous):

Let x = the number of five-dollar bills Since the total number of bills is 38, the number of one-dollar bills will be 38-x Then we can write the following equation for the total $ amount: 1*(38-x) + 5x = 86 Solve for x: 38 - x + 5x = 86 4x = 48 x = 12 Therefore, the number of 5-dollar bills is 12 And the number of 1-dollar bills is 38-12 = 26